OpenStudy (anonymous):

Anabel took two tests this semester and her average is 75. If she scores 100 on the rest of her tests, how many more tests does Anabel have to take to raise her average to 90?

OpenStudy (anonymous):

How do I set this up?

jimthompson5910 (jim_thompson5910):

Let S = sum of the first two scores

jimthompson5910 (jim_thompson5910):

So the average of those two scores is S/2 = 75 which means that S = 75*2 S = 150

jimthompson5910 (jim_thompson5910):

Now let x = number of additional tests she has to take

jimthompson5910 (jim_thompson5910):

If she gets 100 on each test, then you're adding 100x more points onto the total sum (if she takes all x tests and gets 100 each time) This gives you 150 + 100x Divide this by 2+x and set this equal to 90 to get (150+100x)/(2+x) = 90 Now solve for x

OpenStudy (anonymous):

x=3?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so she must take 3 more tests and get 100 on each one

jimthompson5910 (jim_thompson5910):

this is of course assuming everything is weighted equally

OpenStudy (anonymous):

How would I be able to remember how to do this type of word problem on a test?

jimthompson5910 (jim_thompson5910):

i would try to remember the formula for averages, which is A = S/n where A is the average, S is the sum and n is the number of things you're averaging

jimthompson5910 (jim_thompson5910):

that's how I started the problem out from there, you use this formula in a slightly altered way to find the number of additional tests needed

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

:)

jimthompson5910 (jim_thompson5910):

np